OpenStudy (anonymous):

2log4(10) – log4(5^2)

OpenStudy (cwrw238):

what do you want to do with this/ express it as a single logarithm?

OpenStudy (anonymous):

Find the value

OpenStudy (anonymous):

I am a little unclear of how to do it. I got a final answer of 1

OpenStudy (cwrw238):

2log4(10) = log4 100 log4 100 - log4 25 = log4 (100/25) = log4 4 = 1

OpenStudy (anonymous):

I am also working on \[\log _{3\sqrt[3]{9}}\]

OpenStudy (cwrw238):

its a matter of applying the laws of logs

OpenStudy (anonymous):

perfect! i got the same answer the one with the radical sign is kinda of tripping me up as well

OpenStudy (cwrw238):

i cant make that one out its cube root of what?

OpenStudy (anonymous):

cubed root of 9

OpenStudy (cwrw238):

and log to the base 3?

OpenStudy (anonymous):

yep log3 3^√9

OpenStudy (cwrw238):

ok that log3 9^(1/3) = !/3 log3 9 = 1/3 log3 3^2 = 2/3 log3 3 = 2/3

OpenStudy (anonymous):

so 3√9 is actually just 9^1/3?

OpenStudy (anonymous):

ohhh okay I get it, cause you move the exponent over to the otehr side of the log like the other one. perfect thanks so much!

OpenStudy (cwrw238):

oh - i might have misinterpreted that is 3 times the square root of 9 or the cube root of 9?

OpenStudy (anonymous):

I think its the cubed root of 9. like the 3 is in small font infront of the square root symbol

OpenStudy (cwrw238):

right yes cube root of 9 is the same as 9^(1/3)

OpenStudy (anonymous):

than you are suppose to make 9 into an exponent like you did with 3^2? why does 1/3 become 2/3 infront of log?

OpenStudy (cwrw238):

log3 3^2 = 2 log3 3 thats one of laws of logs (!/3) log3 3^2 = (1/3 * 2) log3 3 = 2/3 log 3

OpenStudy (anonymous):

ohhh i see cause you muliply 2 by 1/3. That makes alot of sense. Thanks so much for all the help

OpenStudy (cwrw238):

yw